OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (radar):

calculate slope which is:\[\Delta y \over \Delta x\]

OpenStudy (radar):

\[\Delta y=1-10=-9\]\[\Delta x= -2 -1=-3 \]

OpenStudy (radar):

slope m= -9/-3= 3

OpenStudy (radar):

Now develop the equation: 3 = (y-1)/x-(-2) 3x+6=y-1 3x-y=-7

OpenStudy (radar):

That looks like the first one.